df SS Regression 1 Residual 67 55.35 Total 68 183.44 The ANOVA table above is from a...

df SS
Regression 1
Residual 67 55.35
Total 68 183.44

The ANOVA table above is from a simple linear regression analysis relating the percentage alcohol content in diferent brands of beer to the number of kilojoules per 100mL. Determine the coefficient of determination as a percentage, correct to two decimal places.
Answer #1

Coefficient of determination, R^2 =1-( SSR/TSS) = 1-(55.35/183.44) = 0.6983

So, the anova table or the regression is able to capture 69.83 percent of variation through regression